# DER Day-Ahead Offering: A Neural Network Column-and-Constraint Generation Approach

### v1

@@ IV-A Experiment Setup
To validate the effectiveness of the proposed NN-accelerated CCG algorithm, we test it on a 1028-node synthetic distribution network and is employed for comparison against state-of-the-art benchmark approaches. The synthetic distribution system is equipped with various resources in different busses, including rooftop solar PV (10-kW), behind-the-meter battery energy storage (Tesla’s Powerwall model, 11.3 kW/14.5 kWh), and an average load of 1.3 MW (peaking at 1.7 MW). All experiments are executed in Google Colab on a 64-bit Windows 11 workstation with an Intel® Core™ i7-6700HQ CPU running at 2.64 GHz and 24 GB of RAM, using GUROBI 12.0, and gurobi-machinelearning-1.5.5.dev0 is used to embed NNs into MILPs. The basic unit in this work is a fully-connected multilayer perceptron (MLP) with Rectified Linear Unit (ReLU) activation functions. We assess the performance of NN-accelerated CCG using two metrics: (i) optimality gap, computed as the mean absolute error (MSE) with the traditional CCG solution as the reference, and (ii) computational speedup relative to a classical CCG baseline. To test the performance of proposed NN-accelerated CCG algorithm, we benchmark it with two state-of-the-art methods: 1) Gurobi [gurobi], which can directly solve the 2S-ARSO problem through a large MILP; 2) traditional CCG algorithm [zeng_solving_2013].
@@ IV-B Structure and Accuracy of Neural Network
For data training, we sample problem instances, candidate first-stage decisions, and DER uncertainty scenarios to construct feature vectors encoding constraint coefficients and objective weights. Labels are obtained by solving the second-stage recourse with $(\boldsymbol{x},\boldsymbol{\xi})$ fixed. These data collection procedures are independent and parallelizable. In this day-ahead 2S-ARSO problem, we draw 1,0001{,}000 instances, generate $5$ first-stage decisions per instance, and sample $20$ scenarios per decision, producing 100,000100{,}000 labeled examples split into 80,00080{,}000 for training and 20,00020{,}000 for validation. Both embedding NNs employ two-layer MLPs with layer widths $[64,8]$. The value NN has one hidden layer of width $[8]$. We undergo 500 training epochs with validation MSE computed every 10 epochs. Model selection criteria prioritize minimal mean absolute validation error. Empirical findings reveal close alignment between training and validation MSE, with the NN achieving sub-1.3% MSE, indicating strong accuracy performance.
@@ IV-C Effectiveness of NN-Accelerated CCG
A comparative evaluation of Gurobi, classical CCG, and the proposed NN-accelerated CCG on the synthetic 1028-bus system, across varying numbers of sampled day-ahead price trajectories, is summarized in Table I. As shown in Table I, compared with Gurobi-based method, the proposed algorithm yields reasonable optimal objective value with the gap of $0.1\%$ across all instances. This indicates that the trained ReLU-activated NN can well estimate the worse-case and optimal value of the second-stage recourse problem. Regarding the computational performance, under four number of sampled day-ahead price trajectories from Markov process, the proposed algorithm can achieve at least $21.87\%$ speedup compared with Gurobi-based method and achieve $7.04\%$ speedup over traditional CCG algorithm benchmark. This improved computational efficiency lies in the pre-trained value NN surrogates for the second-stage recourse problem. Under the large number of day-ahead price trajectories, such as the 1028 synthetic network with 500 day-ahead price trajectories, NN-accelerated CCG solve the 2S-ARSO within 1,2321{,}232 seconds, whereas the computational time for Gurobi and traditional CCG is much longer. As the number of sampled day-ahead price trajectories increases and instances become more challenging, the NN-accelerated CCG delivers lower runtimes and improved solution quality. Unlike classical CCG, which alternates between a master problem and recourse problem, our approach embeds a surrogate of the recourse in a single MILP and thus removes decomposition overhead. Across all test cases the trade-off between speed and accuracy is favorable: near-optimal solutions are obtained with less computation time, which is valuable in time-critical settings.
@@ V Conclusion
This paper introduces an NN–accelerated CCG algorithm for the day-ahead optimal offering problem for DER aggregators. A tailored ReLU-activated NN, encoded through MILP constraints, serves as a surrogate for the recourse problem. The NN enables the master problem to incorporate Benders cuts and scenario selections while preserving feasibility. On large synthetic distribution networks, the approach can achieve 1 to 2 orders of magnitude speedup over traditional CCG benchmarks, while maintaining solution optimality. These results demonstrate the effectiveness of the proposed NN-accelerated CCG method in substantially improving the tractability of the network-constrained optimal DER offering problems.
Appendix A Finite-Termination of NN-Accelerated CCG

### v2

@@ IV-A Experiment Setup
To validate the effectiveness of the NN-accelerated CCG algorithm, we test it on a 1028-node synthetic distribution network and is employed for comparison against state-of-the-art benchmark approaches. The synthetic distribution system is equipped with various resources in different busses, including rooftop solar PV (10-kW), behind-the-meter battery energy storage (Tesla’s Powerwall model, 11.3 kW/14.5 kWh), and an average load of 1.3 MW (peaking at 1.7 MW). All experiments are executed in Google Colab on a 64-bit Windows 11 workstation with an Intel® Core™ i7-6700HQ CPU running at 2.64 GHz and 24 GB of RAM, using GUROBI 12.0, and gurobi-machinelearning-1.5.5.dev0 is used to embed NNs into MILPs. The basic unit in this work is a fully-connected multilayer perceptron (MLP) with Rectified Linear Unit (ReLU) activation functions. We assess the performance of NN-accelerated CCG using two metrics: (i) optimality gap, computed as the mean absolute error (MSE) with the traditional CCG solution as the reference, and (ii) computational speedup relative to a classical CCG baseline. To test the performance of NN-accelerated CCG algorithm, we benchmark it with two state-of-the-art methods: 1) Gurobi [6], which can directly solve the 2S-ARSO problem through a large MILP; 2) traditional CCG algorithm [14].
@@ IV-B Structure and Accuracy of Neural Network
For data training, we sample problem instances, candidate first-stage decisions, and DER uncertainty scenarios to construct feature vectors encoding constraint coefficients and objective weights. Labels are obtained by solving the second-stage recourse, which is a tractable deterministic MILP as $(\boldsymbol{x},\boldsymbol{\xi})$ are fixed. These data collection procedures are independent and parallelizable. The data collection times, training times, and total times are 562s, 844s, and 1,314s for the 2S-ARSO under 25 price trajectories, respectively. To keep the master and recourse problems tractable, we train small networks that achieve low MSE; therefore, we do not perform elaborate hyperparameter tuning, as further optimization would primarily refine the already strong numerical results. For the overall 2S-ARSO problem, we draw 1,0001{,}000 instances, generate $5$ first-stage decisions per instance, and sample $20$ scenarios per decision, producing 100,000100{,}000 labeled examples split into 80,00080{,}000 for training and 20,00020{,}000 for validation. Both embedding NNs employ two-layer MLPs with layer widths $[64,8]$. The value NN has one hidden layer of width $[8]$. We undergo 500 training epochs with validation MSE computed every 10 epochs. Model selection criteria prioritize minimal mean absolute validation error. Empirical findings reveal a close alignment between the MSE of the training and testing dataset, with the NN achieving sub-1.3% MSE, indicating strong accuracy performance.
@@ IV-C Effectiveness of NN-Accelerated CCG
A comparative evaluation of Gurobi, classical CCG, and the NN-accelerated CCG on the synthetic 1028-bus system, across varying numbers of sampled day-ahead price trajectories, is summarized in Table I. As shown in Table I, compared with the Gurobi-based method, the proposed algorithm yields a reasonable optimal objective value with a gap of $0.1\%$ across all instances. This indicates that the trained ReLU-activated NN can accurately estimate the worst-case and optimal value of the second-stage recourse problem. Regarding computational performance, under four different numbers of sampled day-ahead price trajectories from the Markov process, the proposed method achieves at least $21.87\,\times$ speedup compared to the Gurobi-based method and $7.05\,\times$ speedup over the traditional CCG benchmark. This improved computational efficiency lies in the pre-trained value NN surrogates for the second-stage recourse problem. Under a large number of day-ahead price trajectories, such as in the 1028 synthetic network with 500 day-ahead price trajectories, the NN-accelerated CCG solves the 2S-ARSO within 1,2321{,}232 seconds, whereas the computational times for Gurobi and the traditional CCG are much longer. As the number of sampled day-ahead price trajectories increases and instances become more challenging, the NN-accelerated CCG delivers lower runtime and improved solution quality. Unlike classical CCG, which alternates between a master problem and a recourse problem, our approach embeds a surrogate of the recourse in a single MILP, thus removing decomposition overhead. Across all test cases, the trade-off between speed and accuracy is favorable: near-optimal solutions are obtained with less computation time, which is valuable in time-critical settings.
@@ V Conclusion
This paper introduces an NN–accelerated CCG algorithm for the day-ahead optimal offering problem for DER aggregators. A tailored ReLU-activated NN, encoded through MILP constraints, serves as a surrogate for the recourse problem. The NN enables the master problem to incorporate Benders cuts and scenario selections while preserving feasibility. On large synthetic distribution networks, the approach can achieve 1 to 2 orders of magnitude speedup over traditional CCG benchmarks while maintaining solution optimality. These results demonstrate the effectiveness of the NN-accelerated CCG method in substantially improving the tractability of the network-constrained optimal DER offering problems. Potential future directions include the consideration of topology changes, which can be addressed by augmenting the training set with representative contingencies and encoding contingency-dependent network coefficients as inputs.